# How to Correctly Report LLM-as-a-Judge Evaluations

### v2

@@ 4 Method to Correctly Report LLM-as-a-Judge Evaluations
Uncertainty Quantification via Confidence Interval.
To quantify uncertainty in $\hat{\theta}$, we derive a confidence interval for $\theta$ that incorporates variance contributions from both the test and calibration dataset (lang2014confidence):
@@
$d\tilde{\theta}=2z_{\alpha}^{2}\left(-(1\!-\!\tilde{\theta})\!\cdot\!\frac{\tilde{q}_{0}(1-\tilde{q}_{0})}{\tilde{m}_{0}}+\tilde{\theta}\!\cdot\!\frac{\tilde{q}_{1}(1-\tilde{q}_{1})}{\tilde{m}_{1}}\right)\!.$ (13)
Impact of Sample Sizes on Confidence-Interval Length.
The confidence interval in (9) reflects uncertainty from both the test and calibration datasets through $\tilde{n}$, $\tilde{m}_{0}$, and $\tilde{m}_{1}$. As these sample sizes increase, the terms inside the square root decrease, resulting in a shorter confidence interval for $\theta$. Because LLM-as-a-judge evaluations can be run at scale with minimal cost, the test-set size $n$ can often be made extremely large. In the limit $n\to\infty$, test-set uncertainty vanishes, and the interval length is determined solely by the calibration sample sizes $m_{0}$ and $m_{1}$. This observation enables practitioners to target a desired interval length and determine the minimal calibration budget required to achieve it. In Section 6, we further show that this regime leads to variance advantages over direct human evaluation.
Figure 3 illustrates how the confidence-interval length decreases as the calibration dataset grows. The dashed curves correspond to calibration datasets with equal numbers of ‘incorrect’ and ‘correct’ instances ($m_{0}=m_{1}$), using $\hat{q}_{0}=0.7$, $\hat{q}_{1}=0.9$, and test-set accuracies $\hat{p}\in{0.3,0.5,0.7,0.9}$. For example, the red dashed curve ($\hat{p}=0.3$) indicates that achieving an interval shorter than $0.1$ requires approximately $m=m_{0}+m_{1}\approx 200$ calibration examples. Because the calibration dataset is collected independently and its label composition can sometimes be influenced through sampling strategies, asymmetric label sizes ($m_{0}\neq m_{1}$) are feasible. This flexibility is important, as the two label types typically contribute asymmetrically to the overall uncertainty.
[…]
Adaptive Allocation to Reduce Confidence-Interval Length.
Motivated by this observation, we introduce an adaptive allocation procedure in Algorithm 1. The algorithm first collects a small pilot calibration sample (e.g., $m_{\mathrm{pilot}}=10$ per label type) to obtain preliminary estimates $(\tilde{q}_{0},\tilde{q}_{1})$, from which it computes the empirical error ratio $(1-\tilde{q}_{0})/(1-\tilde{q}_{1})$. Combining this ratio with the naive accuracy $\hat{p}$ from the test set, the algorithm estimates the optimal allocation of $(m_{0},m_{1})$ that approximately minimizes the confidence-interval length in (9). As shown by the solid curves in Figure 3, this adaptive allocation yields shorter intervals under a fixed calibration budget than symmetric allocation. The optimality of this rule is established in the following section.
A Python implementation that computes the bias-adjusted estimator $\hat{\theta}$ in (8) and the confidence interval in (9) is provided in Appendix C.
@@ 5 Theoretical Justification of Our Method
We provide theoretical guarantees for the bias-adjusted estimator and confidence interval introduced in Section 4. We analyze bias properties, derive asymptotic variance and confidence intervals, and establish the optimality of the proposed label allocation rule for the calibration dataset.
5.1 Mitigating Bias on Point Estimator.
We compare the bias-adjusted estimator $\hat{\theta}$ in (8) with the naive estimator $\hat{p}$ in (3). By the law of total probability,
$p$ $=\mathop{\rm Pr}\nolimits(\hat{Z}=1\mid Z=1)\cdot\mathop{\rm Pr}\nolimits(Z=1)$ (14)
$\quad+\big(1-\mathop{\rm Pr}\nolimits(\hat{Z}=0\mid Z=0)\big)\cdot\big(1-\mathop{\rm Pr}\nolimits(Z=1)\big)$ (15)
$=(q_{0}+q_{1}-1)\cdot\theta+(1-q_{0}).$ (16)
Thus, $\mathbb{E}[\hat{p}]=\theta$ for all $\theta$ if and only if the LLM judge is perfect, i.e., $q_{0}=q_{1}=1$; otherwise, $\hat{p}$ is biased. When $q_{0}+q_{1}<2$, the expectation can be rewritten as
$\displaystyle\mathbb{E}[\hat{p}]=p$ $=\theta+(2-q_{0}-q_{1})\Big(\frac{1-q_{0}}{2-q_{0}-q_{1}}-\theta\Big),$ (17)
@@
Bias-adjusted Estimator.
Assuming $q_{0}+q_{1}>1$, inverting the above relation gives the estimator in Eq. (7). Replacing $(p,q_{0},q_{1})$ with their empirical estimates gives the bias-adjusted estimator $\hat{\theta}$ in (8(rogan1978estimating; lang2014confidence). When $q_{0}$ and $q_{1}$ are known, $\hat{\theta}$ is unbiased; when they are estimated from a calibration dataset, residual bias remains.
The following result shows that, even with estimated sensitivity and specificity, the adjusted estimator $\hat{\theta}$ achieves smaller bias than the naive estimator as the calibration dataset grows. All proofs are provided in Appendix B.
@@
Suppose that $m:=2m_{0}=2m_{1}$ and that $q:=q_{0}=q_{1}$ with $0.5<q\leq 1$. For sufficiently large $m\gtrsim 2q/(2q-1)^{2}$, the absolute bias of $\hat{\theta}$ in (2) is always smaller than that of $\hat{p}$ in (3) for all $\theta$.
Even when $\hat{q}_{0}$ and $\hat{q}_{1}$ are estimated from data, the adjusted estimator $\hat{\theta}$ exhibits lower bias than the naive estimator once the calibration dataset is large enough. Moreover, the required calibration size depends on judge reliability: fewer samples suffice when the LLM has high sensitivity and specificity ($q\approx 1$), while substantially more are required as performance approaches chance level ($q\approx 0.5$).
@@ 5.2 Uncertainty Quantification via Confidence Interval
We quantify uncertainty in $\hat{\theta}$ arising from two sources: the test dataset used to estimate $p$ and the calibration dataset used to estimate $q_{0}$ and $q_{1}$. Applying the delta method (dorfman1938note; ver2012invented) gives the asymptotic variance
$\displaystyle\mathrm{Var}\big(\hat{\theta}\big)\!=\!\frac{\frac{\hat{p}(1\!\!-\!\hat{p})}{n}+(1-\hat{\theta})^{2}\!\cdot\!\frac{\hat{q}_{0}(1-\hat{q}_{0})}{m_{0}}+\hat{\theta}^{2}\!\cdot\!\frac{\hat{q}_{1}(1-\hat{q}_{1})}{m_{1}}}{(\hat{q}_{0}+\hat{q}_{1}-1)^{2}},\,\,\,$ (18)
using the binomial variance formulas for $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$. A detailed derivation is provided in Appendix B.1.
Based on this variance, we construct a confidence interval for $\theta$ using the “add two successes and two failures” adjusted Wald approach (de1820theorie; agresti1998approximate; brown2001interval; lang2014confidence). Specifically, we replace $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$ with their adjusted versions $\tilde{p}$, $\tilde{q}_{0}$, and $\tilde{q}_{1}$, as defined in Eq. (12). These adjustments can be interpreted as adding one (or $z_{\alpha}^{2}/2$) success and one (or $z_{\alpha}^{2}/2$) failure to each estimate, improving coverage accuracy for small sample sizes (agresti2000simple). Substituting the estimates gives the confidence interval in Eq. (9).
The adjustment also induces a small shift in the interval center (i.e., $d\tilde{\theta}$ in Eq. (13)) due to the dependence of $\hat{\theta}$ on $\hat{q}_{0}$ and $\hat{q}_{1}$. Its effect on interval length is negligible and therefore ignored in the final approximation, see lang2014confidence for details.
The Optimal Allocation of the Calibration Dataset.
We characterize how to allocate calibration label sizes under a fixed budget to minimize the confidence-interval length. Assume that $\tilde{q}_{0},\tilde{q}_{1}$ are close to one and define the error ratio $\kappa:=(1-\tilde{q}_{0})/(1-\tilde{q}_{1})$.
@@
This result provides a guideline for allocating calibration samples between the two response types. When the LLM is less accurate at identifying ‘incorrect’ responses (i.e., when the error ratio $\kappa$ is large), more calibration samples should be allocated to estimating $\tilde{q}_{0}$. Moreover, when the proportion $\tilde{p}$ of ‘correct’ predictions in the test set is small, the factor $(1/\tilde{p}-1)$ amplifies the contribution of $\tilde{m}_{0}$ to the interval length, again favoring additional calibration samples for ‘incorrect’ responses.
To implement this rule in practice, we use the adaptive procedure in Algorithm 1. The procedure estimates $(\tilde{q}_{0},\tilde{q}_{1})$ from a small pilot calibration sample, computes the error ratio $\hat{\kappa}$, and allocates $(m_{0},m_{1})$ according to the approximate optimal ratio implied by Proposition 5.2.
6 When Are LLM-as-a-Judge Evaluations Preferable to Fully Human Evaluation?
@@
In Section 5, we show that a human-evaluated calibration dataset is essential for reliable LLM-as-a-judge evaluation, as it enables correction of bias arising from judgment errors. However, this requirement raises a natural question. If human annotators are already available to label a set of instances, one could instead directly estimate the target accuracy $\theta$ by evaluating responses using human labels.
@@
More concretely, given a fixed evaluation budget that allows labeling $m$ responses, there are two possible strategies: (i) use the $m$ labeled instances as a calibration dataset to estimate the sensitivity and specificity of an LLM judge and then correct large-scale LLM-based evaluations, or (ii) directly evaluate $m$ test instances with human labels and estimate the target accuracy from these samples alone.
Since a calibration dataset is required for reliable LLM-as-a-judge evaluation, determining which of these two strategies is preferable is a fundamental question. We address this question from a statistical perspective and show that there exist parameter regimes in which, for the same evaluation budget $m$, LLM-as-a-judge evaluation with calibration is statistically more efficient than fully human evaluation, yielding more precise accuracy estimates.
Regimes in Which LLM-as-a-Judge Evaluations Achieve Lower Variance in Estimation.
Any estimator based on a finite sample incurs non-negligible variance. In fully human evaluation, this variance cannot be arbitrarily reduced because human annotation is costly and does not scale.
In contrast, in LLM-as-a-judge evaluations, the test dataset can be scaled at negligible cost. As the test-set size increases, the variance contributed by LLM-based judgments becomes negligible, and the dominant source of uncertainty arises from bias correction using a finite human-labeled calibration dataset. This variance decreases as the LLM judge becomes more reliable, i.e., as its sensitivity $q_{1}$ and specificity $q_{0}$ approach one. Aggregating large-scale LLM judgments with bias correction can yield an estimator with lower variance than one based solely on human evaluation.
Figure 4 visualizes this comparison. The shaded region indicates values of the true accuracy $\theta$ for which the LLM-as-a-judge estimator under our framework achieves lower variance than direct human evaluation. This region is centered around $\theta=1/2$, where evaluation is intrinsically most uncertain, and expands as the LLM judge becomes more accurate. Intuitively, when the correctness of a test instance is ambiguous and the LLM judge is reliable, aggregating many inexpensive LLM judgments with bias correction is more statistically efficient than human annotation alone.
The following proposition formalizes this comparison by characterizing the parameter regimes in which LLM-as-a-judge evaluation with calibration is statistically more efficient than a fully human evaluation.
@@
[…]
Figure 4: Comparison of estimator variances between LLM-as-a-judge evaluation under our method ($\hat{\theta}$ in Eq. (12)) and direct human evaluation ($\hat{\phi}$ in Proposition 6.1). The shaded regions indicate parameter regimes where the LLM-as-a-judge estimator under our method has lower variance than direct human evaluation.
7 Robust Estimation under Distribution Shift
Having established when LLM-as-a-judge evaluation is preferable to human evaluation, we now consider which LLM-as-a-judge estimators should be preferred. To this end, we compare our estimator with existing calibration-based approaches, highlighting differences in the distributional assumptions required for unbiasedness and how these assumptions behave under distribution shift between the test distribution $\mathbb{P}$} and the calibration distribution $\mathbb{Q}$}.

### v3

@@ 4.1 Mitigating Bias in Point Estimator
4.2 Uncertainty Quantification via Confidence Interval
To quantify uncertainty in $\hat{\theta}$, we derive a confidence interval for $\theta$ that incorporates variance contributions from both the test and calibration dataset (Lang and Reiczigel, 2014):
@@ 4.2 Uncertainty Quantification via Confidence Interval
$d\tilde{\theta}=2z_{\alpha}^{2}\left(-(1-\tilde{\theta})\cdot\tfrac{\tilde{q}_{0}(1-\tilde{q}_{0})}{\tilde{m}_{0}}+\tilde{\theta}\cdot\tfrac{\tilde{q}_{1}(1-\tilde{q}_{1})}{\tilde{m}_{1}}\right).$ (12)
The confidence interval in (8) reflects uncertainty from both the test and calibration datasets through $\tilde{n}$, $\tilde{m}_{0}$, and $\tilde{m}_{1}$. As these sample sizes increase, the terms inside the square root decrease, resulting in a shorter confidence interval for $\theta$. Because LLM-as-a-judge evaluations can be run at scale with minimal cost, the test-set size $n$ can often be made extremely large. In the limit $n\to\infty$, test-set uncertainty vanishes, and the interval length is determined solely by the calibration sample sizes $m_{0}$ and $m_{1}$. This observation enables practitioners to target a desired interval length and determine the minimal calibration budget required to achieve it.
Figure 3 illustrates how the confidence-interval length decreases as the calibration dataset grows. The dashed curves correspond to calibration datasets with $m_{0}=m_{1}$. For example, when $\hat{p}=0.3$ is estimated from the test set and $\hat{q}_{0}=0.7$, $\hat{q}_{1}=0.9$ (red dashed curve), achieving an interval shorter than $0.1$ requires $m\approx 200$ calibration examples.
Figure 3: Confidence-interval length versus calibration-set size under $\hat{q}_{0}=0.7$, $\hat{q}_{1}=0.9$, and $n\to\infty$. We consider four test-set cases in which the naive estimator $\hat{p}$ takes one of the values in $\{0.3,0.5,0.7,0.9\}$. Dashed curves correspond to calibration sets with symmetric allocation across label types ($m_{0}=m_{1}$), while solid curves correspond to calibration sets using the adaptive allocation strategy in Algorithm 1, which results in shorter intervals.
Allocation Strategy to Reduce Confidence-Interval Length.
Furthermore, as the calibration dataset is collected independently, its label composition can sometimes be influenced through sampling strategies. As a result, asymmetric label sizes ($m_{0}\neq m_{1}$) are feasible. This flexibility matters because the two label types typically contribute asymmetrically to the overall uncertainty, so an imbalanced allocation can reduce the interval length.
Motivated by the benefits of asymmetric allocation, we introduce an adaptive allocation strategy in Algorithm 1. The algorithm first collects a small pilot calibration set (e.g., $m_{\mathrm{pilot}}=10$ per label type) to obtain preliminary estimates $\tilde{q}_{0}$ and $\tilde{q}_{1}$. It then computes the empirical error ratio $(1-\tilde{q}_{0})/(1-\tilde{q}_{1})$ and combines it with the test-set estimate $\hat{p}$ to determine an allocation $(m_{0},m_{1})$ that approximately minimizes the confidence-interval length in (8). As shown by the solid curves in Figure 3, this adaptive allocation gives shorter intervals under a fixed calibration budget than symmetric allocation. The optimality of this strategy in minimizing the confidence-interval length under a fixed calibration budget is established in the following section.
@@ 5 Theoretical Justification of Our Method
We provide theoretical justification for the method proposed in Section 4, including derivations, bias analysis, and optimality of the allocation strategy.
5.1 Mitigating Bias in Point Estimator
We compare the bias-corrected estimator $\hat{\theta}$ in (7) with the naive estimator $\hat{p}$ in (3). By the law of total probability,
$\displaystyle\mathbb{E}[\hat{p}]=p=(q_{0}+q_{1}-1)\cdot\theta+(1-q_{0}).$ (13)
Thus, $\mathbb{E}[\hat{p}]=\theta$ for all $\theta$ if and only if the LLM judge is perfect (i.e., $q_{0}=q_{1}=1$); otherwise, $\hat{p}$ is biased. When $q_{0}+q_{1}<2$, the expectation can be rewritten as
@@ 5.1 Mitigating Bias in Point Estimator
Bias-Corrected Estimator.
Assuming $q_{0}+q_{1}>1$, inverting the above relation in (13) gives the estimator in (6). Replacing $q_{0}$ and $q_{1}$ with their empirical estimates $\hat{q}_{0}$ and $\hat{q}_{1}$ gives the bias-corrected estimator $\hat{\theta}$ in (7).
When $q_{0}$ and $q_{1}$ are known, $\hat{\theta}$ is unbiased. When they are estimated from a calibration dataset, $\hat{\theta}$ can exhibit bias. Nevertheless, the following result shows that $\hat{\theta}$ attains smaller bias than the naive estimator $\hat{p}$ in (3) as the calibration-set size grows. All proofs are provided in Appendix B.
@@
Suppose that $m:=2m_{0}=2m_{1}$ and that $q:=q_{0}=q_{1}$ with $0.5<q\leq 1$. For sufficiently large $m\gtrsim 2q/(2q-1)^{2}$, the absolute bias of $\hat{\theta}$ in (2) is always smaller than that of $\hat{p}$ in (3) for all $\theta\in[0,1]$.
Even when $\hat{q}_{0}$ and $\hat{q}_{1}$ are estimated from data, $\hat{\theta}$ has smaller bias than $\hat{p}$ for sufficiently large calibration sets, with the required size depending on the LLM judge’s sensitivity and specificity: fewer samples suffice when the LLM has high sensitivity and specificity ($q\approx 1$), while substantially more are required as these approach chance level ($q\approx 0.5$).
@@ 5.2 Uncertainty Quantification via Confidence Interval
We quantify uncertainty in $\hat{\theta}$ arising from two sources: the test dataset used to estimate $p$ and the calibration dataset used to estimate $q_{0}$ and $q_{1}$. Applying the delta method and the binomial variance formulas for $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$, we obtain the asymptotic variance
$\displaystyle\mathrm{Var}\big(\hat{\theta}\big)\!=\!\frac{\frac{\hat{p}(1\!\!-\!\hat{p})}{n}+(1\!-\!\hat{\theta})^{2}\!\cdot\!\frac{\hat{q}_{0}(1-\hat{q}_{0})}{m_{0}}+\hat{\theta}^{2}\!\cdot\!\frac{\hat{q}_{1}(1-\hat{q}_{1})}{m_{1}}}{(\hat{q}_{0}+\hat{q}_{1}-1)^{2}}.\;\;\;\;\;$ (15)
A detailed derivation is provided in Appendix B.1.
Based on this variance, we construct a confidence interval for $\theta$ using the “add two successes and two failures” adjusted Wald approach (de Laplace, 1820; Agresti and Coull, 1998; Brown et al., 2001; Lang and Reiczigel, 2014). Specifically, we replace $\hat{p}$, $\hat{q}_{0}$, and $\hat{q}_{1}$ with their adjusted versions $\tilde{p}$, $\tilde{q}_{0}$, and $\tilde{q}_{1}$, as defined in (11). These adjustments can be interpreted as adding one (or $z_{\alpha}^{2}/2$) success and one (or $z_{\alpha}^{2}/2$) failure to each estimate, improving coverage accuracy for small sample sizes (Agresti and Caffo, 2000). Substituting the estimates gives the confidence interval in (8).
The adjustment also induces a small shift in the interval center (i.e., $d\tilde{\theta}$ in (12)) due to the dependence of $\hat{\theta}$ on $\hat{q}_{0}$ and $\hat{q}_{1}$. Its effect on interval length is negligible and therefore ignored, see Lang and Reiczigel (2014) for details.
Optimal Allocation for Minimizing Confidence-Interval Length.
We characterize the optimal allocation of calibration samples across label types under a fixed budget for minimizing the confidence-interval length. To this end, we define the error ratio $\kappa:=(1-\tilde{q}_{0})/(1-\tilde{q}_{1})$.
@@
This result provides guidance for allocating calibration samples between the two label types. In particular, when the LLM judge is less accurate at identifying ‘correct’ responses (i.e., when $\kappa$ is small) or when the naive estimator $\tilde{p}$, the proportion of test-set responses judged as ‘correct’ by the LLM, is large, more calibration samples should be allocated to estimating $\tilde{q}_{1}$ to minimize the interval length, resulting in an optimal allocation with $m_{0}<m_{1}$. In practice, we approximate the optimal allocation using Algorithm 1, which estimates the error ratio $\hat{\kappa}$ from a pilot calibration sample and allocates $m_{0}$ and $m_{1}$ according to Proposition 5.2.
6 When Are LLM-as-a-Judge Evaluations Preferable to Human-Only Evaluation?
@@
In Section 5, we show that a human-labeled calibration dataset is essential for reliable LLM-as-a-judge evaluation because it enables bias correction. This raises a natural question: If human annotators are available, why not instead estimate the target accuracy $\theta$ directly from the human-only evaluation on the test set?
@@
More concretely, suppose that a fixed evaluation budget of $m$ human annotations is available. One could either (i) use these $m$ annotations to construct a calibration dataset for estimating the LLM-judge’s sensitivity and specificity, and then apply bias correction to LLM-based evaluations, or (ii) apply them directly to $m$ test instances to estimate the true accuracy from human evaluation alone. We compare these two strategies and show that, for the same budget $m$, LLM-as-a-judge evaluation with our bias-correction method can be statistically more efficient than human-only evaluation in certain parameter regimes.
Parameter Regimes Where LLM-as-a-Judge Achieves Lower Variance.
Human-only evaluation relies on a limited annotation budget, which restricts how many test instances can be evaluated and therefore limits how much the variance of the estimator can be reduced. In contrast, LLM-as-a-judge evaluation can be applied to an arbitrarily large test set at negligible marginal cost, so that the dominant source of uncertainty stems from bias correction based on a finite calibration set. These differences reflect an inherent trade-off between a costly, near-perfect oracle that can be queried only on a small number of instances, versus an inexpensive but imperfect oracle that can be deployed at scale. Importantly, as the judge becomes more reliable (i.e., as $q_{0}$ and $q_{1}$ approach one), the calibration-induced uncertainty diminishes. Consequently, by aggregating large-scale LLM judgments with appropriate bias correction, it is possible to obtain an estimator whose variance is smaller than that of human-only evaluation.
It is worth clarifying why human-only evaluation can still exhibit variance even if the true label of every test instance is evaluated and therefore known. The quantity of interest is the true accuracy of the test distribution, i.e., the expected value of the true label indicating whether an instance is ‘correct’. In practice, we observe a random sample of $m$ instances from the test distribution, and different samples of size $m$ result in different empirical accuracies. Thus, the variance of the estimator arises from subsampling the test distribution, not from uncertainty in individual labels.
Figure 4: Comparison of variances between the LLM-as-a-judge estimator under our correction method, $\hat{\theta}$ in (11), and the human-only estimator, $\hat{\phi}$ in Proposition 6.1, when $m\to\infty$. Shaded regions indicate regimes where our LLM-as-a-judge evaluation is preferable to human-only evaluation in terms of estimator variance.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 5: Monte Carlo simulation for estimating $\theta$ under an imperfect LLM judge with $(q_{0},q_{1})=(0.7,0.9)$. We evaluate estimators across 21 values of $\theta\in[0,1]$, each visualized as a single point. Figure 5(a) reports the results from a single run, while Figure 5(b) and Figure 5(c) summarize averages computed over 10,000 replications. All experiments use a test dataset of size $n=1000$ and a calibration dataset of size $m=200$, and we use an equal allocation $m_{0}=m_{1}$ for Figure 5(a) and Figure 5(b). (a) The naive estimator $\hat{p}$ in (3) exhibits bias, while the bias-corrected estimator $\hat{\theta}$ in (7) closely recovers the true accuracy $\theta$ across all values. Shaded regions represent the 95% confidence intervals (CI). (b) Across all $\theta$, the coverage probability of the confidence interval remains consistently close to the nominal 95% level. (c) Given a fixed calibration budget of $m=200$, we compare two allocation strategies: an equal split ($m_{0}=m_{1}$) and the allocation proportional to $m_{0}\propto(1/\hat{p}-1)\sqrt{\kappa}\cdot m_{1}$ by using Algorithm 1. The proposed allocation gives shorter confidence intervals.
Figure 4 visualizes this comparison. The shaded regions indicate parameter regimes over $q_{0}$, $q_{1}$, and $\theta$ where the bias-corrected LLM-as-a-judge estimator has smaller variance than the human-only estimator. The variance advantage is most pronounced around $\theta=1/2$, where the intrinsic uncertainty of binary evaluation is largest, and the shaded region expands as the LLM judge becomes more accurate. The following proposition formalizes these parameter regimes.
@@
$\!\!\!\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}\!\cdot\!\left(1+\frac{(1-\theta)\epsilon}{1-\theta-\epsilon}+\frac{\theta\epsilon}{\theta-\epsilon}\right)\!.\;$ (16)
Moreover, since $\epsilon=O(m^{-1/2})$, as $m\to\infty$ the sufficient condition in (16) reduces to the necessary and sufficient condition $\theta(1-\theta)\geq\frac{q(1-q)}{(2q-1)^{2}}$, which is equivalent to
$\theta\in\left[\tfrac{1}{2}-\sqrt{\tfrac{1}{2}-\tfrac{1}{4(2q-1)^{2}}},\tfrac{1}{2}+\sqrt{\tfrac{1}{2}-\tfrac{1}{4(2q-1)^{2}}}\right].$ (17)
7 Empirical Validation of Our Method
We validate the theoretical results in Section 5 through Monte Carlo simulations and real-world benchmarks.
7.1 Monte Carlo Simulation
We evaluate the proposed method under the following parameter configuration. The LLM judge is characterized by $(q_{0},q_{1})=(0.7,0.9)$, and the true accuracy varies over $\theta\in\{0,0.05,0.10,\ldots,1\}$, yielding 21 settings. For each $(q_{0},q_{1},\theta)$, we generate a test dataset of size $n=1000$ and a calibration dataset of total size $m=200$, with equal allocation $m_{0}=m_{1}$ unless stated otherwise.
We report the naive estimator $\hat{p}$ in (3) and its confidence interval, as well as the bias-corrected estimator $\hat{\theta}$ in (7) with the confidence interval in (8). Each configuration is replicated 10,000 times to evaluate estimation accuracy, confidence-interval coverage, and interval length. Additional results are provided in Appendix D.
Results.
Figure 5(a) compares $\hat{p}$ and $\hat{\theta}$ in a single simulation run. The naive estimator $\hat{p}$ is biased, particularly overestimating the true accuracy $\theta$ when $\theta$ is small. In contrast, the bias-corrected estimator $\hat{\theta}$ closely matches the true accuracy across all values of $\theta$, consistent with Proposition 5.1.
Figure 5(b) reports the empirical coverage probability of the confidence interval based on $\hat{p}$ and the proposed confidence interval in (8). The interval based on $\hat{p}$ attains near-zero coverage except at a few values of $\theta$. In contrast, the proposed interval achieves coverage close to the nominal 95% level across all values of $\theta$. These results confirm that our confidence interval provides reliable uncertainty quantification.
To evaluate the benefit of optimal allocation, we compare symmetric allocation ($m_{0}=m_{1}=100$) with the allocation produced by Algorithm 1 under a fixed calibration budget of $m=200$. Algorithm 1 produces a calibration set that approximately follows the optimal allocation in Proposition 5.2. Figure 5(c) shows that Algorithm 1 yields shorter confidence intervals than symmetric allocation.
7.2 Chatbot Arena Benchmark
We evaluate our method on the Chatbot Arena benchmark (Chiang et al., 2024), where users vote on which of two anonymous models provides the better response to the same prompt. These pairwise comparisons are aggregated into each model’s win rate, defined as the probability that the target model’s response is preferred over its opponent’s.
We consider six target models studied in Zheng et al. (2023): Alpaca-13B (Taori et al., 2023), Claude-v1, FastChat-T5-3B, GPT-4, LLaMA-13B (Touvron et al., 2023), and Vicuna-13B (Chiang et al., 2023). For each target model, we construct response pairs consisting of one response generated by the target model and one response generated by a randomly selected opponent model. Chatbot Arena provides human preference labels for these pairs, which we treat as ground truth for computing the target model’s true win rate $\theta$. We then apply GPT-4.1-mini as an LLM judge to the same pairs to obtain LLM-judged preference labels, and compute the naive estimator $\hat{p}$ in (3) and our bias-corrected estimator $\hat{\theta}$ in (7).
For each trial, we randomly split the human-labeled pairs into a test set (90%) and a calibration set (10%). We estimate the judge’s sensitivity and specificity on the calibration set, and evaluate the bias of $\hat{p}$ and $\hat{\theta}$ on the held-out test set. We repeat this procedure 100 times with independent random splits and report results averaged across trials.
Figure 6: Average bias of winning rates for six models on Chatbot Arena, averaged over 100 random test (90%) / calibration (10%) splits. We compare the naive LLM-based estimator $\hat{p}$ in (3) with our bias-corrected estimator $\hat{\theta}$ in (7), using GPT-4.1-mini as the judge. The proposed method reduces bias for all models.
Table 1: Empirical coverage of the 95% confidence intervals in (8) and average interval lengths for Chatbot Arena win-rate estimates.
Model Alpaca Claude FastChat-T5 GPT-4 LLaMA Vicuna
Coverage 96% 95% 97% 97% 99% 99%
CI length 0.193 0.262 0.256 0.210 0.267 0.178